# Physical modelling of global macrosystems evolution

### v1

@@ Abstract
The dynamics of growth and innovation often exhibit sudden, explosive surges, where systems remain quasi stable for extended periods before accelerating dramatically—often surpassing traditional exponential growth. This pattern is evident across various domains, including world population increases and rapid technological advancements. Although these phenomena share common characteristics, they are driven by different underlying mechanisms. In this paper, we introduce a unified framework to capture these phenomenologies through a theory of combinatorial innovation. Inspired by the Theory of the Adjacent Possible, we model growth and innovation as emerging from the recombination processes of existing elements of a system. By formalizing these qualitative ideas, we provide a mathematical structure that explains diverse phenomena, enables cross-system comparisons, and offers grounded predictions for future growth trajectories. Our approach distils the complexity of innovation into a more accessible yet robust framework, paving the way for a deeper and more flexible mathematical understanding of growth and innovation processes.
@@ I Introduction
Many systems—ranging from biological Markov and Korotayev (2007); Grinin et al. (2015); Panov (2004) and social Von Foerster et al. (1960); Kapitza (1996); Korotayev (2013) to economic LePoire (2013); Korotayev (2005)—share a striking pattern of explosive growth and innovation. When we look at metrics like world population size Kremer (1993), average GDP per capita Korotayev (2009), or historical milestones Kurzweil (2004); LePoire (2015), the pace of increase over time is undeniable. These systems often grow at rates that surpass traditional exponential models Korotayev and LePoire (2020); Steffen et al. (2015), instead following hyperbolic functions Korotayev (2006); Panov (2005). This brings us to the concept of the ’singularity,’ where growth theoretically approaches infinity in a finite time. In real-world systems, this singularity can be interpreted as a harbinger of profound transformations or disruptive events on the horizon. While hyperbolic models fit the data impressively Korotayev (2009); Panov (2020), their true power and ability to explain real-world phenomena remain elusive. Additionally, the microscopic mechanisms driving such rapid growth are poorly understood, often leading to applying the same mathematical models to systems with very different behaviors.
@@
The theory of ”singularity” is primarily phenomenological. Many innovation-driven systems have been effectively interpreted through combinatorial processes Koppl et al. (2018); Cortês et al. (2022a). In this framework, innovation and growth emerge from recombining elements within the system Kauffman (2000); Koppl et al. (2023), similar to assembling existing parts to create new tools. The Theory of the Adjacent Possible (TAP) Cortês et al. (2022b) proposes that systems can expand into their ”possible” space, which consists of elements that are not yet realized but are just one step away from being created  Kauffman (2000). Models based on this concept, such as the TAP equation, have been used to describe the growth of the world population Koppl et al. (2018) and GDP Koppl et al. (2023). Moreover, the concept of the Adjacent Possible has also been applied within the framework of Urn Models Tria et al. (2014); Di Bona et al. (2023), which reproduce key features of novelties dynamics, such as Zipf’s, Heaps’ and Taylor’s laws Tria et al. (2018); Loreto et al. (2016).
@@
The combinatorial approach effectively replicates empirical data on a qualitative level Koppl et al. (2018); Steel et al. (2020) but often lacks a more quantitative and mathematical analysis. In this work, we aim to reconcile the hyperbolic growth observed in natural systems and the phenomenological theory of ”singularity” with the microscopically grounded theory of combinatorial innovation. This will allow us to distinguish between seemingly similar systems yet governed by different underlying processes. We demonstrate the versatility of our framework by applying it to a wide range of systems, including world population Korotayev (2006); Kapitza (1996), biological and cosmological phase shifts Kurzweil (2005); Korotayev (2020); Panov (2004), GDP per capita Korotayev (2005), and US patent records Youn et al. (2015).
@@
We will end by discussing the concept of ”singularity”, which is often interpreted as a point where the extreme growth rate would lead to a qualitative change in the system’s behavior Kurzweil (2004); Panov et al. (2020); Korotaev (2006); Modis (2013); LePoire (2020). We argue that in real-world systems, which are discrete and finite, such a concept must be carefully considered, particularly when making predictions.
@@ II The TAP equation: a paradigm for combinatorial growth
Figure 1: Evolution of Canonical Milestones and Biological Phase Transitions over time. Left. Macrodevelopment rate $\rho$ as a function of time, computed as the inverse of the inter-event time between successive milestones. The log-linear scale highlights the hyperbolic growth pattern, as exponential growth would appear as a straight line on this scale. Solid lines show the results of the fit discussed in Appendix C. Center. The same plot on a log-log scale, where the hyperbolic growth manifests as a straight line. The power-law exponent is $-1$, indicating a relationship of the form $y=1/x$. Right. Number of events as a function of time, plotted on a linear-log scale to emphasize the logarithmic behavior of milestone accumulation over time. The solid line represents equation 4, using the same parameters as those fitted to the macrodevelopment rate $\rho$.
@@
The idea that innovations and growth arise as a combinatorial process is closely connected to the Theory of the Adjacent Possible (TAP) proposed by Kauffman et al. Kauffman (2000); Cortês et al. (2022b). The concept underlying the TAP is that the range of potential future developments is generated by the existing elements within the system, creating a space of possibilities where novel elements emerge from combinations of pre-existing ones. This framework applies to both physical entities (e.g., molecules, genes, tools) and conceptual ones (e.g., patents, songs, ideas).
@@
This space of possibilities, known as the Adjacent Possible, contains all the unexplored elements that are just one step away from realization Kauffman (2000). Mathematically, this concept led to simple yet powerful models known as the TAP equation. If we denote by $M_{t}$ the number of elements in the system at time $t$, their evolution is governed by the number of combinations that can be formed from these elements:
$M_{t+1}=(1-\mu)M_{t}+\sum_{i=1}^{M_{t}}\alpha_{i}\binom{M_{t}}{i}$ (1)
@@
Here, $\alpha_{i}$ represents coefficients that account for the realization probability of each combination, with typically decreasing values reflecting the increasing difficulty of forming longer combinations. For simplicity, we assume no extinction, setting $\mu=0$. We consider all terms in the summation starting from $i=1$, though some versions of the TAP equation start from $i=2$. These models are known for their explosive behavior, with growth rates exceeding exponential ones. TAP models have effectively described the growth of various systems, such as world population Koppl et al. (2018), GDP per capita, and the emergence of new technologies Koppl et al. (2023).
Analyzing equation 1 reveals that its analytic continuation diverges in finite time Cortês et al. (2022b). Hyperbolic functions naturally describe this behavior, featuring a singularity where the function diverges. However, it is important to recognize that this mathematical concept doesn’t apply to the real world, as no physical discrete system can exhibit infinite growth in finite time. Despite this limitation, hyperbolic functions have been widely used to interpret various systems, such as world population Korotayev (2006) and biological phase shifts in the field of biocosmology Korotayev (2020).
While the TAP equation offers a robust physical framework, its connection to hyperbolic functions has often been overlooked. The analytical continuation of the TAP equation naturally leads to hyperbolic expressions. For instance, in the simplest version of a TAP equation, with no extinction rate ($\mu=0$) and uniform coefficients ($\alpha_{i}=\alpha$ for any $i$), the model simplifies to:
$M_{t+1}=M_{t}+\alpha\sum_{i=1}^{M_{t}}\binom{M_{t}}{i}$
@@
Using Pascal’s Triangle to evaluate the combinatorial term and considering the analytical continuation, we obtain:
$\frac{dM_{t}}{dt}=\alpha(2^{M_{t}}-1)\approx\alpha 2^{M_{t}}$
@@
The solution to this differential equation is a hyperbolic function:
$M_{t}=\log_{2}\left(\dfrac{1}{2^{-M_{0}}-\alpha(t-t_{0})}\right)$ (2)
@@
where $M_{0}$ is the initial number of elements in the system at time $t_{0}$. This is a hyperbolic expression featuring a singularity at time $t=t_{0}+\alpha^{-1}2^{-M_{0}}$. This result illustrates how hyperbolic trends observed in real-world data can arise from combinatorial models. In the following sections, we will demonstrate how different TAP models replicate various behaviors in real-world scenarios.
@@ III.1 Biological Phase Transitions
In the realm of biocosmology, numerous studies have explored the emergence of paradigmatic events that have significantly influenced the history of the Universe and Life. These events include the origin of the Milky Way, the advent of life, the Cambrian explosion, the Hominoid revolution, and the emergence of democracy. Some studies refer to them as Canonical Milestones Kurzweil (2004, 2005); Modis (2003), while others as Biological Phase Transitions Panov (2004); Panov et al. (2020). Despite slight variations in data representation, both terms refer to the same phenomena with similar meanings and intentions. The two corresponding time series are presented in Appendix A.
When examining the time series for the occurrence of these milestones, authors from both perspectives have recognized a hyperbolic growth behavior (see Figure 1). Instead of analyzing the direct count of milestones, most researchers consider the ”inter time” Panov (2004); Korotayev (2020), which measures the time between two consecutive milestones. The inverse of the inter time, which represents the macrodevelopment rate $\rho_{t}$ and is effectively an approximation of the derivative of the milestone count $dM_{t}/dt$, closely follows a hyperbolic function with an exponent of $-1$:
$CM\ :\ \rho_{t}=\frac{2.054}{t-2029},\quad PT\ :\ \rho_{t}=\frac{1.886}{t-2027}$
for Canonical Milestones (CM) and Biological Phase Transitions (PT), respectively.
Starting with the general TAP equation 1 and assuming no extinction rate ($\mu=0$), our choice of the model relies on the selection of the parameter form for $\alpha$. Given that we know no specific mechanisms leading to the emergence of these phase transitions, we assume that any possible recombination of existing elements could be significant. To account for the increasing difficulty of combining a greater number of elements, we follow the idea presented in Cortês et al. (2022b) and set the coefficients to scale as a power law, $\alpha_{i}=\alpha^{i}$ for $\alpha\in(0,1)$. With this choice, the continuous TAP equation can be reformulated as Cortês et al. (2022b):
$\frac{dM_{t}}{dt}=\sum_{i=1}^{M_{t}}\alpha^{i}\binom{M_{t}}{i}\approx(1+\alpha)^{M_{t}}=e^{\beta M_{t}}$ (3)
with $\beta=\ln(1+\alpha)$. The solution to this equation leads to the following hyperbolic expression (details are reported in Appendix B):
$M_{t}\approx\frac{1}{\beta}\log\left(\dfrac{1}{\beta(t-t_{0})+e^{-\beta M_{0}}}\right)$ (4)
where $M_{0}$ is the initial condition of the system at time $t_{0}$. By deriving the number of the milestones $M_{t}$, we can compute the macrodevelopment rate $\rho_{t}$, which follows a hyperbolic function with an exponent of $-1$:
$\rho_{t}=\frac{dM_{t}}{dt}\approx\dfrac{1}{\beta(t-t_{0})+e^{-\beta M_{0}}}$ (5)
This model replicates the same analytical behavior observed in previous analyses Panov (2004); Kurzweil (2005). Importantly, these expressions emerge naturally from the combinatorial model’s structure, offering a straightforward physical interpretation rather than being derived solely from observed time series data. We can fit the time series using our model and compute the coefficient $\alpha$ that reproduces the actual number of milestones $M_{t}$. The results of this numerical analysis are presented in Figure 1 and in Appendix C.
The values of the coefficients in the two cases are $\alpha_{CM}=1.16\cdot 10^{-10}$ and $\alpha_{PT}=1.58\cdot 10^{-10}$. These model parameters represent the fraction of combinations occurring at any given time $t$, which indicates the rate of successful combinations that actually take place at each time step. We can refer to $\alpha$ as the macrodevelopment parameter of the system. The small difference in their values for Canonical Milestones and Biological Phase Transitions is due to the different events considered in the two time-series. However, the two parameters are comparable, reflecting a similar underlying process.
Additionally, the expressions we derived present a singularity at $t=t_{0}+\beta^{-1}e^{-\beta M_{0}}$, a time value at which the function theoretically diverges. However, the discrete version of the model in equation 3 does not exhibit any divergence in finite time, as it happens for real-world systems. We will explore the significance of this concept in Section IV.
@@
Figure 2: World population and GDP growth over time. Left. World population size from 1000 to 2020 CE. The sudden explosion in the last century suggests a rate that exceeds exponential growth. The solid line represents the hyperbolic fit, as discussed in Appendix C. Center. GDP aggregated for all countries, in millions of dollars, from 700 to 2018 CE. The hyperbolic growth pattern is again evident from the sharp rise in GDP in recent years. The solid line shows the hyperbolic fit, as detailed in Appendix C. Right World population and GDP evolution on a log-log scale. This scale highlights the hyperbolic growth, with a power law exponent of $1$ for world population and $2$ for GDP.

### v2

# Modelling Singularities in Macroevolution

@@ Abstract
Macroevolutionary dynamics often display sudden, explosive surges, where systems remain relatively stable for extended periods before experiencing dramatic acceleration that frequently exceeds traditional exponential growth. This pattern is evident in biological evolution, cultural shifts, and technological progress and is often referred to as the emergence of singularities. Despite their widespread occurrence, these explosions arise from distinct underlying mechanisms in different domains. In this context, we present a unified framework that captures these dynamics through a theory of combinatorial innovation. Building on the Theory of the Adjacent Possible, we model macroevolutionary change as a process driven by recombining pre-existing elements within a system. By formalising these qualitative insights, we provide a mathematical structure that explains the emergence of these explosive phenomena, facilitates comparisons across different systems, and enables predictive insights into future evolutionary trajectories. Moreover, by comparing discrete and continuous formalisations of the theory, we emphasise that the occurrence and observation of these presumed singularities should be carefully considered, as they arise from the continuous limit of inherently discrete models.
@@ I Introduction
Many systems—ranging from biological markov2007phanerozoic ; grinin2015modeling ; panov2004avtomodel and social von1960doomsday ; kapitza1996phenomenological ; korotayev2013globalization to economic lepoire2013potential ; korotayev2005compact —share a striking pattern of explosive growth and innovation. When we look at metrics like world population size kremer1993population , average GDP per capita korotayev2009compact , or historical milestones kurzweil2004law ; lepoire2015interpreting , the pace of increase over time is undeniable. These systems often grow at rates that surpass traditional exponential models korotayev202021st ; steffen2015trajectory , instead following hyperbolic functions korotayev2006world ; panov2005scaling . This brings us to the concept of the ’singularity,’ where growth theoretically approaches infinity in a finite time. In real-world systems, this singularity can be interpreted as a harbinger of profound transformations or disruptive events on the horizon. While hyperbolic models fit the data impressively korotayev2009compact ; panov2020singularity , their true power and ability to explain real-world phenomena remain elusive. Additionally, the microscopic mechanisms driving such rapid growth are poorly understood, often leading to applying the same mathematical models to systems with very different behaviours.
@@
The theory of ”singularity” is primarily phenomenological. Many innovation-driven systems have been effectively interpreted through combinatorial processes koppl2018simple ; cortes2022biocosmology . In this framework, innovation and growth emerge from recombining elements within the system kauffman2000investigations ; koppl2023explaining , similar to assembling existing parts to create new tools. The Theory of the Adjacent Possible (TAP) cortes2022tap proposes that systems can expand into their ”possible” space, which consists of elements that are not yet realised but are just one step away from being created  kauffman2000investigations . Models based on this concept, such as the TAP equation, have been used to describe the growth of the world population koppl2018simple , GDP koppl2023explaining , and technological innovations sole2016singularities . Moreover, the concept of the Adjacent Possible has also been applied within the framework of Urn Models tria2014dynamics ; di2025dynamics , which reproduce key features of novelties dynamics, such as Zipf’s, Heaps’ and Taylor’s laws tria2018zipf ; loreto2016dynamics .
@@
The combinatorial approach effectively replicates empirical data on a qualitative level koppl2018simple ; steel2020dynamics but often lacks a more quantitative and mathematical analysis. In this work, we aim to reconcile the hyperbolic growth observed in natural systems and the phenomenological theory of ”singularity” with the microscopically grounded theory of combinatorial innovation. This allows us to distinguish between seemingly similar systems yet governed by different underlying processes. We demonstrate the versatility of our framework by applying it to a wide range of scenarios, including world population korotayev2006world ; kapitza1996phenomenological , biological and cosmological phase shifts kurzweil2005singularity ; korotayev2020twenty ; panov2004avtomodel , GDP per capita korotayev2005compact , and US patent records youn2015invention .
@@
The modelisation we propose also sheds light on the concept of ”singularity”, which is often interpreted as a point where the extreme growth rate would lead to a qualitative change in the system’s behaviour kurzweil2004law ; panov2020twenty ; korotaev2006introduction ; modis2013singularity ; lepoire2020exploring . We argue that in real-world systems, which are discrete and finite, such a concept must be carefully considered, particularly when making predictions.
@@ II The TAP equation: a paradigm for combinatorial growth
The idea that innovations and growth arise as a combinatorial process is closely connected to the Theory of the Adjacent Possible (TAP) proposed by Kauffman et al. kauffman2000investigations ; cortes2022tap . The concept underlying the TAP is that the range of potential future developments is generated by the existing elements within the system, creating a space of possibilities where novel elements emerge from combinations of pre-existing ones. This framework applies to both physical entities (e.g., molecules, genes, tools) and conceptual ones (e.g., patents, songs, ideas).
@@
This space of possibilities, known as the Adjacent Possible, contains all the unexplored elements that are just one step away from realisation kauffman2000investigations . Mathematically, this concept led to simple yet powerful models known as the TAP equations. If we denote by $M_{t}$ the number of elements in the system at time $t$, their evolution is governed by the number of combinations that can be formed from these elements:
$M_{t+1}=(1-\Sigma)M_{t}+\sum_{i=1}^{M_{t}}A_{i}\binom{M_{t}}{i}$ (1)
@@
Here, $A_{i}$ represents the coefficients that account for the probability of realisation of each combination, with typically decreasing values reflecting the increasing difficulty of forming longer combinations. For simplicity, we assume no extinction, setting $\Sigma=0$. We consider all terms in the summation starting from $i=1$, although some versions of the TAP equation start from $i=2$. These models are known for their explosive behaviour, with growth rates exceeding exponential ones. Models based on the TAP concept have been used to describe the growth of the world population koppl2018simple and GDP koppl2023explaining . A similar approach has been explored by Solé et al. sole2016singularities , who proposed a generalised model of technological evolution. Their work incorporates ageing effects to study how recombinatorial dynamics influence innovation rates and prevent singularities in some cases. These examples emphasize the versatility of combinatorial frameworks to describe technological and innovation-driven systems.
While the TAP equations describe the evolution of a (temporal) discrete system, most real world situations are characterised by a continuous time. Therefore we need to derive the continuous limit of (1) in order to apply the TAP equations to a continuous scenario. For the passage to the continuum, we pose $t\rightarrow t_{k}=\frac{k}{n}T=\delta k$ for $\ T\gg 1$ and $\delta=T/n$ so that $\lim_{n\rightarrow\infty,T\rightarrow\infty}\delta=0$. In addition, we pose: $\Sigma=\delta\sigma$ and $A_{i}=\delta\overline{\nu}\alpha_{i}$. For simplicity, we introduce the dimensional factor $\overline{\nu}$, which has the dimension of a rate, ensuring that the model parameters $\alpha_{i}$ are dimensionless.
In this way, one has:
$M_{t_{k+1}}=(1-\Sigma)M_{t_{k}}+\sum_{i=1}^{M_{t_{k}}}A_{i}\binom{M_{t_{k}}}{i},$
from which we get:
$\frac{M_{t_{k+1}}-M_{t_{k}}}{\delta}=-\frac{\Sigma}{\delta}M_{t_{k}}+\sum_{i=1}^{M_{t_{k}}}\frac{A_{i}}{\delta}\binom{M_{t_{k}}}{i},$
and, in the limit $\delta\rightarrow 0$, one obtains the continuous limit of Eq. ((1)) as:
$\frac{dM_{t}}{dt}=-\sigma M_{t}+\overline{\nu}\sum_{i=1}^{M_{t}}\alpha_{i}\binom{M_{t}}{i},$ (2)
Note that $\sigma$ has the dimension of a rate, while the $\alpha_{i}$ are dimensionless due to the explicit introduction of the dimensional factor $\overline{\nu}$.
Previous works already noticed that in the continuous limit, Eq. ((1)) leads to divergences in finite times cortes2022tap . Despite this, the direct relation between the TAP equations, hyperbolic functions and the emergence of a singularity has been previously overlooked. Indeed, the continuous limit of the TAP equations naturally leads to hyperbolic expressions. For instance, we can consider the simplest case, with no extinction rate ($\Sigma=0$) and uniform coefficients ($A_{i}=A$ for any $i$). In this way (1) simplifies to:
$M_{t+1}=M_{t}+A\sum_{i=1}^{M_{t}}\binom{M_{t}}{i}$ (3)
@@
Using Pascal’s Triangle to evaluate the combinatorial term and taking the limit of continuous time, we obtain, analogously to (2):
$\frac{dM_{t}}{dt}=\overline{\nu}\alpha(2^{M_{t}}-1)\approx\overline{\nu}\alpha 2^{M_{t}}.$ (4)
@@
The solution to this differential equation is the following function with a logarithmic divergence:
$M_{t}=\log_{2}\left(\dfrac{1}{2^{-M_{0}}-\overline{\nu}\alpha\ln 2(t-t_{0})}\right)$ (5)
@@
where $M_{0}$ is the initial number of elements in the system at time $t_{0}$. The singularity of such equation lies at time $t^{*}=t_{0}+(\overline{\nu}\alpha\ln 2)^{-1}2^{-M_{0}}$. This result illustrates how singular trends observed in real-world data can arise from combinatorial models.
Before diving into the detailed analysis of real-world systems, it is worth providing some remarks about the concept of singularity and its definition. Mathematically speaking, a function $f(t)$ has a singularity in $t^{*}$ if $t^{*}$ is an accumulation point of $f$. This occurs when the function diverges as $t$ approaches a value $t^{*}<\infty$:
$\lim_{t\to t^{*}}f(t)=\infty,\qquad t^{*}<\infty.$
It is important to stress that discrete models, despite a clear explosive behaviour, only have an accumulation point in $\infty$. This means that $f(t)$ can diverge only as $t\to\infty$. In other words, while we can define the discrete model also for $t>t^{*}$, i.e., we can compute $M_{t}$ for any $t<\infty$, in the continuous approximation the function $M_{t}$ does not exist for times larger than the singularity, so for $t>t^{*}$. Despite this limitation, hyperbolic functions have been widely used to interpret various systems, such as world population korotayev2006world and biological phase shifts in the field of biocosmology korotayev2020twenty .
Figure 1: Evolution of Canonical Milestones (CM) and Biological Phase Transitions (BT) over time. Left. Macrodevelopment rate $\rho_{t}$ as a function of time, computed as the inverse of the inter-event time between successive milestones. The log-linear scale highlights the hyperbolic growth pattern, as exponential growth would appear as a straight line on this scale. Solid lines show the results of the fit discussed in Appendix C. Time is expressed as the number of years before 2020 CE. Center. The same plot on a log-log scale, where the hyperbolic growth manifests as a straight line. The power-law exponent is $-1$, indicating a relationship of the form $y=1/x$. Right. Number of events $M_{t}$ as a function of time, plotted on a linear-log scale to emphasize the logarithmic behaviour of milestone accumulation over time. The solid line represents the fit from the model in Eq. (6).
@@ III.1 Biological Phase Transitions
Figure 2: World population and GDP growth over time. Left. World population size from 1000 to 2020 CE. The sudden explosion in the last century suggests a rate that exceeds exponential growth. The solid line represents the hyperbolic fit, as discussed in Appendix C. Time is expressed as the number of years before 2020 CE. Center. GDP aggregated for all countries, in billions of USD dollars (indicated as GDP and adjusted for Parity for Purchase Power bolt2020maddison ), from 700 to 2018 CE. The hyperbolic growth pattern is again evident from the sharp rise in GDP in recent years. The solid line shows the hyperbolic fit, as detailed in Appendix C. Right World population and GDP evolution on a log-log scale. This scale highlights the hyperbolic growth, with a power law exponent of $1$ for world population and $2$ for GDP.
In the realm of biocosmology, numerous studies have explored the emergence of paradigmatic events that have significantly influenced the history of the Universe and Life. These events include the origin of the Milky Way, the advent of life, the Cambrian explosion, the Hominoid revolution, and the emergence of democracy. Some studies refer to them as Canonical Milestones kurzweil2004law ; kurzweil2005singularity ; modis2003limits , while others as Biological Phase Transitions panov2004avtomodel ; panov2020twenty . Despite slight variations in data representation, both terms refer to the same phenomena with similar meanings and intentions. The two corresponding time series are presented in Appendix A.
When examining the time series for the occurrence of these milestones, authors from both perspectives have recognised a hyperbolic growth behaviour (see Figure 1). Instead of analysing the direct count of milestones, most researchers consider the ”inter time” panov2004avtomodel ; korotayev2020twenty , which measures the time between two consecutive milestones. The inverse of the inter time, which represents the macrodevelopment rate $\rho_{t}$ and is effectively an approximation of the derivative of the milestone count $dM_{t}/dt$, closely follows a hyperbolic function with an exponent of $-1$:
$CM\ :\ \rho_{t}=\frac{2.054}{2029-t},\quad BT\ :\ \rho_{t}=\frac{1.886}{2027-t}$
for Canonical Milestones (CM) and Biological Phase Transitions (BT), respectively. Here, time $t$ represents the number of years in the Common Era (CE), meaning that years before 0 CE (commonly referred to as BCE, Before Common Era) are represented as negative values.
Starting with the general TAP equation (1) and assuming no extinction rate ($\mu=0$), our choice of the model relies on the selection of the parameter form for $\alpha$. Given that we know no specific mechanisms leading to the emergence of these phase transitions, we assume that any possible recombination of existing elements could be significant. To account for the increasing difficulty of combining a greater number of elements, we follow the idea presented in koppl2018simple ; cortes2022tap and set the coefficients to scale as a power law, $\alpha_{i}=\alpha^{i}$ for $\alpha\in(0,1)$. With this choice, the continuous TAP equation can be reformulated as cortes2022tap :
$\frac{dM_{t}}{dt}=\overline{\nu}\sum_{i=1}^{M_{t}}\alpha^{i}\binom{M_{t}}{i}\approx\overline{\nu}(1+\alpha)^{M_{t}}=\overline{\nu}e^{\beta M_{t}}$ (6)
with $\beta=\ln(1+\alpha)$, and $\overline{\nu}$ is the standard dimensional factor accounting for the time scale koppl2018simple ; cortes2022tap . Its value can be directly inferred from the data, as explained in Appendix C.
From this expression, we derive the evolution of the macrodevelopment rate $\rho_{t}=\frac{dM_{t}}{dt}$:
$\dfrac{d\rho_{t}}{dt}=\dfrac{d}{dt}\left(\dfrac{dM_{t}}{dt}\right)=\beta\dfrac{dM_{t}}{dt}\overline{\nu}e^{\beta M_{t}}\approx\beta\rho_{t}^{2}$ (7)
which matches the dynamics proposed in previous studies korotayev202021st . The scaling factor $\overline{\nu}$ disappears in the equation for the rate, as any rescaling of time proportionally affects both the rate and the time variable. The solution of this differential equation is a hyperbolic function with an exponent of $-1$:
$\rho_{t}\approx\dfrac{1}{\rho_{0}^{-1}-\beta(t-t_{0})}$ (8)
The explicit expression for the number of milestones $M_{t}$ can be obtained from equation (6) (details are provided in Appendix B):
$M_{t}\approx\frac{1}{\beta}\ln\left(\dfrac{1}{e^{-\beta M_{0}}-\overline{\nu}\beta(t-t_{0})}\right)$ (9)
This model replicates the same analytical behaviour observed in previous analyses panov2004avtomodel ; kurzweil2005singularity . Importantly, these expressions emerge naturally from the combinatorial model’s structure, offering a straightforward physical interpretation rather than being derived solely from observed time series data. We can fit the time series using our model and compute the coefficient $\alpha$ that reproduces the actual number of milestones $M_{t}$. The results of this numerical analysis are presented in Figure 1 and in Appendix C. Following a convention widely adopted in the literature korotayev2020twenty , the time axis is shifted to represent the number of years before the present. Specifically, we use the horizontal axis $\tau=2020-t$, where $2020$ CE is taken as the reference year.
The values of the coefficients in the two cases are $\alpha_{CM}=0.670$ and $\alpha_{BT}=0.837$. These model parameters represent the fraction of combinations occurring at any given time $t$, which indicates the rate of successful combinations that actually take place at each time step. We can refer to $\alpha$ as the macrodevelopment parameter of the system. The small difference in their values for Canonical Milestones and Biological Phase Transitions is due to the different events considered in the two time-series. However, the two parameters are comparable, suggesting a similar underlying process. It is also important to note that these parameter values cannot be used for long-term predictions of the time series. This limitation arises from the increasing divergence between the continuous model and the discrete one as the system approaches the singularity.
Additionally, the expressions we derived present a singularity at $t=t_{0}+\beta^{-1}\rho_{0}^{-1}$, a time value at which the function theoretically diverges. However, the discrete version of the model in equation (6) does not exhibit any divergence in finite time, as it happens for real-world systems. We will explore the significance of this concept in Section IV.